OpenStudy (mrhirohito):

What is the value of the expression? A. –22 B. –14 C. 14 D. 30

OpenStudy (mrhirohito):

@griefer101

OpenStudy (griefer101):

a negative times a negative will be a positive so(-6)x(-3)=?

OpenStudy (mrhirohito):

18?

OpenStudy (griefer101):

yes now subtract 4

OpenStudy (mrhirohito):

14

OpenStudy (griefer101):

yes good job

OpenStudy (mrhirohito):

-14 or 14? xd

OpenStudy (griefer101):

14

OpenStudy (griefer101):

C

OpenStudy (mathstudent55):

Remember the order of operations. You have an addition and a multiplication. Multiplication comes before addition, that's why you need to multiply (-6) * (-3) first. Also, negative tomes negative is positive. -4 + (-6) * (-3) = = -4 + 18 Now you do the addition. = 14

OpenStudy (griefer101):

exactly

OpenStudy (mathstudent55):

@griefer101 Great explanation!

OpenStudy (griefer101):

yep because your not supposed to give the direct answer